OpenStudy (anonymous):

I am not sure what the answer is but I wanna make sure that my answer is correct. The description below represents Function A and the table represents Function B: Function A The function is 5 more than 3 times x. Function B x y -1 2 0 5 1 8

OpenStudy (anonymous):

Which statement is correct about the slope and y-intercept of the two functions? Their slopes are equal, but y-intercepts are not equal. Their slopes are not equal, but y-intercepts are equal. Both slopes and y intercepts are equal. I think it's this one Neither slopes nor y intercepts are equal.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I understand that the equation for function A is written like this y = 5 + 3x

OpenStudy (anonymous):

btw, i'm not saying you're right or wrong... just want to know your thinking

OpenStudy (anonymous):

very good

OpenStudy (anonymous):

and for B?

OpenStudy (anonymous):

Well, heres what I did -1 * 3 = -3 -3 + 5 = 2

OpenStudy (anonymous):

then 0 * 3 = 0 0 + 5 = 5

OpenStudy (anonymous):

1 * 3 = 3 3 + 5 = 8

OpenStudy (anonymous):

very good! you can see from the table data that every time you increase x by 1, y increases by 3. that's the slope. and when x = 0, y = 5, that's the y intercept. great job and you are correct!

OpenStudy (anonymous):

okay thanks! :0

OpenStudy (anonymous):

wait sorry that :0 was supposed to be a :) lol

OpenStudy (anonymous):

no worries!